Solve the problem as directed. Use proportions to find the solution.
The gravitational attraction between two masses varies inversely as the square of the distance between them. If the attraction force between them is 75 lbs. when the bodies are eight feet apart, find the attraction when the masses are twelve feet apart.

Found 2 solutions by stanbon, dfrazzetto:
Answer by stanbon(75887)

The gravitational attraction between two masses varies inversely as the square of the distance between them. If the attraction force between them is 75 lbs. when the bodies are eight feet apart, find the attraction when the masses are twelve feet apart.
----
a = k/d^2
Solve for "k" using "the attraction force between them is 75 lbs. when the bodies are eight feet apart"
75 = k/8^2
k = 4800
-----------------
Equation:
a = 4800/d^2
Find the attraction when the masses are twelve feet apart.
---
a = 4800/12^2
a = 33 1/3 lbs.
=====================
Cheers,
Stan H.

Answer by dfrazzetto(283)


Attraction = 1/(distance)^2, so
75 = 1 / 8^2
75 = 1/64 just as A = 1 / 12^2, so
75*64 = 144A
A = 75*64/144
A = 4800/144 = 33 1/3 lbs √